OpenStudy (maddy1251):

Write the expression 5In M+ 3In N -2IN P - In R as a single logarithm. If someone could start for me or walk me through it, I'd appreciate it..

OpenStudy (anonymous):

Start with these rules: \[a \ln b=\ln b^a\] \[\ln (ab)=\ln a + \ln b\] \[\ln \frac{ a }{ b }=\ln a - \ln b\] Basically the coefficients become exponents on the part after the corresponding log. When you combine added logs go in the numerator, and subtracted ones go in the denominator. An example: \[3 \ln x + 4 \ln y - 5 \ln z\] \[=\ln x^3 + \ln y^4 - \ln z^5\] \[=\ln \frac{ x^3y^4 }{ z^5 }\]

OpenStudy (maddy1251):

\[\ln \frac{ m^{5} n^{3} }{ P^{2} R}\]

OpenStudy (maddy1251):

Is that the answer @peachpi ?

OpenStudy (anonymous):

yes :)

OpenStudy (maddy1251):

@peachpi thank you you're a life saver!! :)

OpenStudy (anonymous):

yw